Given the circle with center O, radius OA, diameter BC, where OA perpendicular BC, and OA=4cm find the length of AB?
Answer by Edwin McCravy(20055)




Since OA is 4 and is a radius, and since OB is also a radius,
OB = 4 too. The Pythagorean theorem says

AB² = OA² + OB²
AB² = 4² + 4²
AB² = 16 + 16
AB² = 32
AB = Ö32 = Ö(16*2) = 4Ö2 or about 5.657 cm.

Edwin